A booster station compresses 50 kg/sec of methane from 14 atm, 285 K to 92 atm in a reversible and isothermal steady state steady-flow manner. Use the general ized fugacity coefficient chart (Item B 37) and calculate the compression work and the heat transfer. Now solve using ideal gas theory and compare answers.

1
Expert's answer
2021-12-13T17:06:04-0500

"W=nRTln\frac{P2}{P1}"

"n=\frac{16.04}{50}"

"=0.3208"

R=518.28

"W=0.3208*518.28*285*\frac{92}{14}"

=311389.14J

=311.39kJ

Heat transfer=work

Heat transfer=311.39kJ